OpenStudy (anonymous):

How would you solve 2x^2-4x=-3? I just want to know if I'm right so far and how to simplify it further.

OpenStudy (studygurl14):

have you completed the square?

OpenStudy (anonymous):

So far I have \[(4\pm4i \sqrt{2})\div4\]

OpenStudy (anonymous):

Not completely sure if I did it right...

OpenStudy (noelgreco):

I get 2i sqrt 2

OpenStudy (noelgreco):

Instead of 4i sqrt 2

OpenStudy (anonymous):

Is that right though? I was unsure if I was supposed to divide by the 4...and it wouldn't it be \[1\pm2i \sqrt{2}\]...?

OpenStudy (anonymous):

*and wouldn't it be

OpenStudy (anonymous):

WAIT! Noooo, I see where I messed up! Thanks n_n

OpenStudy (noelgreco):

\[\frac{ 2\pm i \sqrt{2} }{ 2 }\] Got it?

OpenStudy (anonymous):

Yes! Thank you! I made a dumb mistake before and was getting a wayy wrong answer.